THE FAR-INFRARED SPECTRUM OF DEUTERIUM IODIDE (DI)

\underline{THOMAS D. VARBERG}, JOHN C. ROBERTS, and KAI A. TUOMINEN, Department of Chemistry, Macalester College, St. Paul, MN 55105; KENNETH M. EVENSON, Time and Frequency Division, National Institute of Standards and Technology, Boulder, CO 80303.

Pure rotational transitions within the ground state of deuterium iodide (DI) have been measured over the range J′′ = 2 to 10 by tunable far-infared (TuFIR) spectroscopy. The spectra show large electric-quadrupole splittings arising from the ^{127}I nuclear spin (I = 5/2). A least-squares fit to the measured transitions has resulted in accurate values for the molecular parameters B, D, eQq, D_{eQq} and C_{I}.